Age Problem
1. Jeff is 10 yrs.old and his brother is 2 yrs.old . In how many yrs.will Jeff be twice as old as his brother?

Answer by fractalier(6550)

Call x the number of years in the future.
Then we have
10 + x = 2(2 + x)
10 + x = 4 + 2x
Subtract 4 and subtract x from both sides and get
x = 6 years
(My name is Jeff, but my kid brother is 13 years younger than me!)